Two ships leave port at the same time. One sails south at 15mph and the other sails east at 2o mph. Find a function that models the distance (D) between the ships in terms of the time (T) in hours elasped since thier departure.
I have solved in the followinf way and don't understand where I've gone wrong. I'm coming up with a different answer than what's in my textbook.
(By Pathagorean Therom)
D= 15(t) squared + 20(t) squared.
D(T) = 35(t) squared.
My textbook says that the answer is D(T) = 25T.
How are they getting this solution? Where am I going wrong? Please help!

Answer by stanbon(57290)

Two ships leave port at the same time. One sails south at 15mph and the other sails east at 2o mph. Find a function that models the distance (D) between the ships in terms of the time (T) in hours elasped since thier departure.
--------------------
Draw the picture.
East distance = 20t
South distance = 15t
-------------------------------
Using Pythagoras:
D^2 = (20t)^2 + (15t)^2
---
D^2 = 400t^2 + 225t^2
---
D^2 = 625t^2
----
D = 25t
==========================
Cheers,
Stan H.